# mCLM: A Modular Chemical Language Model that Generates Functional and Makeable Molecules

### v2

@@ 3.3 Library Design: Identifying Functionally-Important Blocks
As a by-product, mCLM is also able to identify blocks which are preferred for certain functions. This can help inform virtual screening campaign design (e.g., which 10 blocks should we use to get the best chemical space for BBBP?), and it can also be useful for stimulating scientific inquiry.
@@
As an example, we calculated the most frequent modifications preferred by the mCLM to improve DILI in FDA drugs (Figure 4). Notably, modification 3 resembles a newly discovered modification for reducing amphotericin B toxicity by human scientists in Maji et al. (2023).
There are many new drug candidates that almost reach FDA approval but fall short for various reasons when tested in clinical trials. For example, Evobrutinib is a Bruton’s tyrosine kinsase (BTK) inhibitor that went through clinical trial as a drug for relapsing Multiple Sclerosis. However, the FDA placed a partial clinical hold on Phase III trials in April 2023 after two patients showed signs of drug-induced liver injury (Montalban et al., 2024). TNG348 is a USP1 inhibitor designed for treating BRCA1/2-mutant and HRD cancers, but it failed in phase 1/2 clinical trials due to liver abnormalities (Inc, 2024; Simoneau et al., 2025). These “fallen angels” represent tremendous opportunities for impactful engagement of the AI/chemistry interface, because much is known about the strengths of each of these small molecules, and it is also known why they fell short. Fixing such fallen angels is a high leverage opportunity for the function-infused mCLM to contribute.
Figure 5: Examples of fallen angel property modification.
Figure 5 shows an application of the mCLM, without synthesis restrictions on the vocabulary, to these two molecules. For both, the initial step is to optimize DILI, the reason the drugs failed in clinical trials. Following that, the mCLM fixes other properties which were made worse in the previous attempt (PGP for Evobrutinib and BBBP for TNG248). For good measure, another property of each molecule is then improved. Notably, at each step, the mCLM only makes minor modifications to each drug of roughly 1 building block. While the mCLM shows promising results for repairing these drugs, it is worth noting that drug discovery is a many-objective optimization problem. While we are able to generate molecules with improved toxicity relative to Evobrutinib and TNG348, as well as other properties, yet other important properties may still have been compromised. Future work may want to investigate longer reasoning chains across a wider variety of properties. For comparison, we show a (less-successful) version of this experiment using MoleculeSTM A.3.
@@ 4 Related Work
Inspired from the successes of protein LLMs, LLMs have been increasingly applied in computational chemistry (Zhang et al., 2024a; b; Fang et al., 2023; Livne et al., 2024; Pei et al., 2023; Zhao et al., 2024; 2023; Yu et al., 2024; Ye et al., 2025; Liu et al., 2023b; Li et al., 2024c; Liu et al., 2024a; Taylor et al., 2022). Recent work has further explored cross-modal modeling of molecule and language (Edwards et al., 2022b; Liu et al., 2022; Li et al., 2025b; Chang and Ye, 2024). General-purpose LLMs such as GPT-4o (OpenAI, 2025a), GPT-5 (OpenAI, 2025b), Gemini-2.5-Flash (Google DeepMind, 2025), and Claude-3.5-H (Anthropic, 2025) show multimodal reasoning capabilities. However, most molecular language models still rely on atom-level (Wang et al., 2019; Ahmad et al., 2022; Zhou et al., 2023; Xia et al., 2023) or Byte-Pair Encoding on SMILES (Chang and Ye, 2024), and thus they often produce invalid tokens, misalign with chemical structures, and generate unsynthesizable molecules. To address these limitations, other approaches incorporate functional groups and fragments into molecular representations (Li et al., 2023a; Nguyen et al., 2024c; Han et al., 2023; Jin et al., 2024; Nguyen et al., 2024a; Wang et al., 2023; Zhang et al., 2020; 2021), capturing both atomic- and group-level information. However, these methods tend to break molecules at bonds that are difficult or even impossible to form through chemical reactions. In contrast, our mCLM decomposes small molecules into function-infused and synthesis-friendly building blocks and integrates synthesis constraints early in training, akin to the way peptide/protein LLMs tokenize at the amino acid level.
@@ 5 Conclusions and Future Work
We have developed a function- and synthesis-aware modular Chemical-Language Model (mCLM), as the first attempt to jointly model natural language sequences with modular chemical language. By design, the mCLM only generates chemical building blocks that can be iteratively assembled on robotic small molecule synthesis platforms, enabling the rapid creation of novel molecules with desired functions, all accessible by non-specialists. In the future we aim to incorporate richer multimodal knowledge related to physical and chemical properties from 2D/3D molecular structures, protein, cell lines and nucleic acid sequences to further enable the mCLM to reason on biological activity, protein docking knowledge, and individuals’ genetic profiles. We plan to extend chemical reasoning to additional aspects such as filling in unspoken knowledge gaps, thinking outside of the box, System 2 thinking for counterfactual reasoning and plausibility prediction, and resolving conflicting claims. We also plan to leverage more physical constraints from simulation tools and chemical and reaction knowledge bases. In the long term, we envision the mCLM as part of a comprehensive, multi-agent, human-in-the-loop autonomous laboratory, structured around iterative cycles of reasoning, proposal, synthesis, physical testing, feedback, and reasoning to enable never-ending self-improvement and co-evolvement with human scientists.
@@ D.2.3 Ensemble Oracle Model:
To generate high-quality synthetic labels for downstream tasks, we construct oracle models focused on ADMET property prediction. We select tasks from the Therapeutics Data Commons (TDC) benchmark  [^3] https://tdcommons.ai/benchmark/admet_group/overview/ using the following criteria:
@@ F.1 Synthesis-Guaranteed Tokenizer Details
The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a pre-determined small set of reactions, preferably only those that can be performed in an automated manner. For instance, if amide bond formation is defined as available, the tokenizer will be able to disconnect amide bonds in the molecule of interest. Up to this point, the protocol is synonymous with classical computational retrosynthesis, but there is a fundamental difference. Namely, the sets of reactions suitable for automated synthesis is very limited – in fact, state-of-the-art synthesis machines utilize only three types of disconnections (amide bond formation, as well as Suzuki and Buchwald-Hartwig couplings). This places very stringent requirements on the groups that can be present in the disconnected blocks – for instance, when the disconnection (say, Buchwald-Hartwig coupling) yields an amine functionality on one of the blocks, this block cannot contain any groups that during the anticipated uses of this block would present a synthetic incompatibility. In the most trivial case, the block cannot contain another unprotected amine because after the Buchwald-Hartwig disconnection, the block would feature two amines which, in turn, would present competing reactive sites (in the Buchwald-Hartwig synthesis but also in the formation of amide bonds). Therefore, the tokenizer performs retrosynthetic operations while simultaneously checking if they do not lead to blocks with functional groups presenting competing reactivities. Only disconnections avoiding such problems are allowed. This then guarantees that when the corresponding blocks are used to make other molecules, they give only the selective synthesis outcomes.

### v3

@@ 3.4 Library Design: Identifying Functionally-Important Blocks
As a by-product, mCLM is also able to identify blocks which are preferred for certain functions. This can help inform virtual screening campaign design (e.g., which 10 blocks should we use to get the best chemical space for BBBP?), and it can also be useful for stimulating scientific inquiry. As an example, we calculated the most frequent modifications preferred by the mCLM to improve DILI in FDA drugs (Figure 5). Notably, modification 3 resembles a newly discovered modification for reducing amphotericin B toxicity by human scientists in Maji et al. (2023).
@@ 4 Related Work
Inspired from the successes of protein LLMs, LLMs have been increasingly applied in computational chemistry (Zhang et al., 2024a; b; Fang et al., 2023; Livne et al., 2024; Pei et al., 2023; Zhao et al., 2024; 2023; Yu et al., 2024; Ye et al., 2025; Liu et al., 2023b; Li et al., 2024c; Liu et al., 2024a; Taylor et al., 2022). Recent work has further explored cross-modal modeling of molecule and language (Edwards et al., 2022b; Liu et al., 2022; Li et al., 2025b; Chang and Ye, 2024; Liu et al., 2024c). General-purpose LLMs such as GPT-4o (OpenAI, 2025a), GPT-5 (OpenAI, 2025b), Gemini-2.5-Flash (Google DeepMind, 2025), and Claude-3.5-H (Anthropic, 2025) show multimodal reasoning capabilities. However, most molecular language models still rely on atom-level (Wang et al., 2019; Ahmad et al., 2022; Zhou et al., 2023; Xia et al., 2023) or Byte-Pair Encoding on SMILES (Chang and Ye, 2024), and thus they often produce invalid tokens, misalign with chemical structures, and generate unsynthesizable molecules. To address these limitations, other approaches incorporate functional groups and fragments into molecular representations (Li et al., 2023a; Nguyen et al., 2024c; Han et al., 2023; Jin et al., 2024; Nguyen et al., 2024a; Wang et al., 2023; Zhang et al., 2020; 2021), capturing both atomic- and group-level information. However, these methods tend to break molecules at bonds that are difficult or even impossible to form through chemical reactions. In contrast, our mCLM decomposes small molecules into function-infused and synthesis-friendly building blocks and integrates synthesis constraints early in training, akin to the way peptide/protein LLMs tokenize at the amino acid level.
Several recent papers propose chemically meaningful tokenization schemes. SAFE (Noutahi et al., 2024) modifies SMILES representations to avoid fragment splits but still inherits core issues of line notations as we described in Section 1, including validity and synthesis challenges. GroupSELFIES (Cheng et al., 2023) extends SELFIES with group-level tokens and can in principle represent our synthesis-friendly blocks. If it is used in a line notation form, however, standard LLM tokenizers will not ensure validity or synthesizability. Reasyn (Lee et al., 2025) is contemporary to our work. It focuses on synthesizable molecule projection using reinforcement learning, requiring the model to predict reactions between pairs of blocks. Unlike Reasyn, mCLM does not require the prediction of reactions, since one is guaranteed to exist between any blocks. Their task and training methodology may be relevant for future mCLM models. We also note recent efforts in learning discrete molecule representations with VQ-like mechanisms (Guo et al., 2025; Ha et al., 2025; Boget et al., 2024). However, these approaches do not guarantee synthesis compatibility and generally lack alignment with natural language. In Section H, we include comparisons with DGAE and HierVAE as additional baselines.
@@ 5 Conclusions and Future Work
We have developed a function- and synthesis-aware modular Chemical-Language Model (mCLM), as the first attempt to jointly model natural language sequences with modular chemical language. By design, the mCLM only generates chemical building blocks that can be iteratively assembled on robotic small molecule synthesis platforms, enabling the rapid creation of novel molecules with desired functions, all accessible by non-specialists. In the future we aim to scale mCLM to larger backbones, incorporate richer multimodal knowledge related to physical and chemical properties from 2D/3D molecular structures, protein-ligand complexes, cell lines and nucleic acid sequences to further enable the mCLM to reason on biological activity, protein docking knowledge, and individuals’ genetic profiles. We plan to extend chemical reasoning to additional aspects such as filling in unspoken knowledge gaps, thinking outside of the box, System 2 thinking for counterfactual reasoning and plausibility prediction, and resolving conflicting claims. We also plan to leverage more physical constraints from simulation tools and chemical and reaction knowledge bases. In the long term, we envision the mCLM as part of a comprehensive, multi-agent, human-in-the-loop autonomous laboratory, structured around iterative cycles of reasoning, proposal, synthesis, physical testing, feedback, and reasoning to enable never-ending self-improvement and co-evolvement with human scientists.
@@
Acknowledgments
We would like to thank Kyunghyun Cho, Anna Hart, Hyeonjeong Ha, Gabriele Scalia, Yanru Qu, and Jeonghwan Kim for helpful discussion. This work is based upon work supported by U.S. DARPA ECOLE Program No. #HR00112390060, and NSF Molecule Maker Lab Institute, an AI Institute for Molecular Discovery, Synthesis Strategy, and Manufacturing funded by the U.S. National Science Foundation under Awards No. 2019897 and 2505932, and NSF NAIRR award. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. This research used the Delta and DeltaAI advanced computing and data resources, which are supported by the National Science Foundation (award OAC 2320345 and award OAC 2005572) and the State of Illinois. Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications. We would like to acknowledge NVIDIA Corporation’s contributions to this work through a grant of NVIDIA A100 Tensor Core GPUs.
@@ D.2.3 Ensemble Oracle Model:
To generate high-quality synthetic labels for downstream tasks, we construct oracle models focused on ADMET property prediction. The oracle models we use for ADMET assessment are state-of-the-art, high-performing models trained on experimental benchmark datasets with strong reported correlation to assay measurements, as shown in Table LABEL:ensemble below. This setup is highly meaningful in computational drug discovery (especially compared to rougher and less meaningful QED scores) and is designed to emulate digital screening pipelines used in industry.
We select tasks from the Therapeutics Data Commons (TDC) benchmark  [^2] https://tdcommons.ai/benchmark/admet_group/overview/ using the following criteria:
@@ F.1 Synthesis-Guaranteed Tokenizer Details
The synthesis-guaranteed tokenizer disconnects the molecule only at bonds that can be formed by a pre-determined small set of reactions, preferably only those that can be performed in an automated manner. For instance, if amide bond formation is defined as available, the tokenizer will be able to disconnect amide bonds in the molecule of interest. Up to this point, the protocol is synonymous with classical computational retrosynthesis, but there is a fundamental difference. Namely, the sets of reactions suitable for automated synthesis is very limited – in fact, state-of-the-art synthesis machines utilize only three types of disconnections (amide bond formation, as well as Suzuki and Buchwald-Hartwig couplings). This places very stringent requirements on the groups that can be present in the disconnected blocks – for instance, when the disconnection (say, Buchwald-Hartwig coupling) yields an amine functionality on one of the blocks, this block cannot contain any groups that during the anticipated uses of this block would present a synthetic incompatibility. In the most trivial case, the block cannot contain another unprotected amine because after the Buchwald-Hartwig disconnection, the block would feature two amines which, in turn, would present competing reactive sites (in the Buchwald-Hartwig synthesis but also in the formation of amide bonds). Therefore, the tokenizer performs retrosynthetic operations while simultaneously checking if they do not lead to blocks with functional groups presenting competing reactivities. Only disconnections avoiding such problems are allowed. This then guarantees that when the corresponding blocks are used to make other molecules, they give only the selective synthesis outcomes. The tokenization process maintains the integrity of the molecules: each molecule can be reconstructed exactly from the building-block sequence, whether produced by tokenization or generated by the mCLM. Each tokenized building block contains placeholder atoms (e.g., [1*], [2*], [3*]) that encode connection points. These placeholders allow us to preserve the full structural integrity of the original molecule, enabling lossless reconstruction from a sequence of tokens. For example, as shown in Figure 1, the drug imatinib with SMILES string
Cc1ccc(cc1Nc2nccc(n2)c3cccnc3)NC(=O)c4ccc(cc4)CN5CCN(CC5)C
is tokenized into:
[ "[3*]c1cccnc1",
  "[2*]c1ccnc(N[1*])n1",
  "[1*]Nc1ccc(C)c([2*])c1",
  "[3*]C(=O)c1ccc(CN2CCN(C)CC2)cc1" ]
Here, [3*] connects to [2*], [1*] connects to [2*], and so on. This process is fully deterministic and reversible. The placeholder atoms and their positions are also encoded by the GNN encoder, allowing the model to reason about both structural and functional context during generation.
Following the connection rules embedded in the placeholders, the original molecule can be fully reconstructed with the following steps:
1. ["[3*]c1cccnc1", "[2*]c1ccnc(N[1*])n1]" → "[1*]Nc1nccc(-c2cccnc2)n1"
   (following the rule [3*] connects to [2*].
2. ["[1*]Nc1nccc(-c2cccnc2)n1", "[1*]Nc1ccc(C)c([2*])c1"]
    → "[1*]Nc1ccc(C)c(Nc2nccc(-c3cccnc3)n2)c1"
   (following the rule [1*] connects to [2*]).
3. ["[1*]Nc1ccc(C)c(Nc2nccc(-c3cccnc3)n2)c1", "[3*]C(=O)c1ccc(CN2CCN(C)CC2)cc1]"
    → "Cc1ccc(NC(=O)c2ccc(CN3CCN(C)CC3)cc2)cc1Nc1nccc(-c2cccnc2)n1"
   (following the rule [1*] connects to [3*]).
@@ Appendix G Source Datasets and Databases
Appendix H Additional Baselines
We conducted additional comparisons with two relevant baselines: DGAE [Boget et al., 2024] (a discrete graph autoencoder using vector quantization) and HierVAE [Jin et al., 2020] (a hierarchical VAE for molecular graph generation). Property scores are in Table 8 and sythesizability scores in Table 9. A unique HierVAE model is trained individually for each property using data from the mCLM corpus, giving it an unfair advantage in per-property optimization, while mCLM operates via instruction prompting. Interestingly, HierVAE performs quite well on synthesizability; we believe that to some extent, this is because HierVAE is trained using molecules constructed from mCLM tokenizer outputs. Thus, it is able to take advantage of the distribution of molecules created by mCLM preprocessing, which likely increases it’s synthesizability performance. Also interestingly, we find a disagreement between the less-rigorous SAscore and actual retrosynthesis metrics, where mCLM outperforms significantly on the more grounded retrosynthesis score. We speculate this is because SAScore prefers simpler molecules or those with simpler substructures, whereas retrosynthesis tools evaluate actual synthetic feasibility rather than only based on structural patterns. However, we note that the difference in synthesizability is still small. Notably, Allchemy may also simply be too strong for the molecules in this distribution, which reduces the dynamic range of the metric. These specific FDA-approved drugs may be too simple, so the dynamic range of molecules predicted synthesizable by Allchemy is quite restricted. For example, even models like MoleculeSTM have > 90% synthesizability when editing these molecules. Future work may need to compare models on more difficult distributions of molecules to increase the dynamic range.
Additionally, we find that HierVAE is effective for optimizing properties (Table 8), especially properties that already have good scores, such as CYP3A4 and HIA. Still, mCLM outperforms it on average.
Table 8: Comparison of pharmacokinetic and toxicity property scores between mCLM and HierVAE. Note that HierVAE is fine-tuned separately for each property, whereas mCLM is used in a instruction-following setting.
Model AMES (↓) BBBP (↑) CYP3A4 (↓) DILI (↓) HIA (↑) PGP (↓) Avg. Improv.
HierVAE 48.2 66.4 1.2 53.1 99.7 64.3 10.5%
mCLM (Ours) 44.4 85.2 1.4 53.7 98.99 64.4 15.0%
Table 9: Synthetic accessibility (SA) [Ertl et al., 2009], validity, and retrosynthetic results for HierVAE. Synthesizability is the percent of valid molecules where a retrosynthetic route was found. Makeability is the overall percent of generations which can be synthesized (Makeability =Valid $\times$ Synth.).
Model SA $(\downarrow)$ Validity (%) Synthesizability (%) Makeability (%)
FDA 2.70 100.0 98.11 98.11
mCLM (Ours) 2.43 100.0 98.23 98.23
HierVAE 2.35 99.70 95.83 95.55
We would also like to explain the technical differences between mCLM and vector-quantized methods, and why we adopt our current approach instead of a vector-quantization technique. Vector-quantized autoencoders (e.g., VQ-VAE) are commonly used to build discrete vocabularies for downstream models. However, in the context of molecular design, these learned codebooks lack chemically grounded guarantees. In contrast, our vocabulary consists of synthesis-verified building blocks, derived from retrosynthesis constraints and domain knowledge, ensuring that all generated molecules are compatible with automation-friendly synthesis. Further, applying vector quantization in this domain risks collapsing chemically distinct structures into the same token, reducing interpretability and limiting extension to new blocks at inference-time. Instead, we use a GNN-based encoder pretrained on chemical properties to produce semantically meaningful embeddings of molecular blocks, which are aligned with a language model via adapter layers. This design allows mCLM to retain both chemical fidelity and generation flexibility, while supporting plug-and-play extensions and instruction-based editing. In Table 10, we compare synthesizability of the DGAE and mCLM on the QM9 dataset. Here, we leverage an existing DGAE model trained on QM9, released by the authors [Boget et al., 2024]. The mCLM is restricted to use synthesis-guaranteed QM9 blocks. In this experiment, the difference is quite striking– the use of our specialized blocks instead of a codebook nearly doubles the synthesizability of generated molecules.
Table 10: Synthesizability comparison using Allchemy on QM9 molecules.
Model Validity (%) Synthesizability (%) Makeability (%) SA Score (↓)
DGAE (VQ-based) 100.0 62.0 62.0 4.46
mCLM (Ours) 100.0 100.0 100.0 2.23
Appendix I Ablation Studies
As detailed in the main text, we performed two ablations of the mCLM block encoding process. First, we trained a SMILES-based version of the mCLM without a GNN (No GNN), still using a Qwen-3B backbone. Second, we conducted a test of the mCLM using blocks extracted from the BRICS algorithm instead of from our tokenization method (No Synth. Tokenizer). These showed significantly degraded property modification capabilities compared to the mCLM.
We also did an ablation study of using Llama-3.2-3B as the backbone for mCLM. We find that, as also noted by others in the community, with the same backbone size, Qwen works better when fine-tuned on downstream tasks than Llama. Specifically, we find evidence that the Llama-3.2-3B-mCLM is overfitting compared with Qwen-2.5-3B-mCLM, evidenced by a lower training loss but higher validation loss (3.702 v.s. 3.620). We also evaluated the property optimization capability of Llama-3.2-3B-mCLM as follows:
Table 11: Comparison of pharmacokinetic and toxicity property scores between using Qwen-2.5-3B and Llama-3.2-3B as backbones.
Model AMES (↓) BBBP (↑) CYP3A4 (↓) DILI (↓) HIA (↑) PGP (↓) Avg. Improv.
mCLM (Llama-3.2-3B) 48.2 59.4 1.7 59.0 98.1 63.9 2.82%
mCLM (Qwen-2.5-3B) 44.4 85.2 1.4 53.7 98.99 64.4 15.0%
mCLM (No GNN) 50.4 46.0 2.5 50.2 97.9 71.7 -7.53%
mCLM (No Synth. Tokenizer) 48.7 55.0 2.9 54.9 98.3 68.3 -19.0%